OpenStudy (anonymous):

Help Final exam test review question!! sketch the region rewrite the integral by switching order evaluate integral

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{y}^{1}x^{2}*e^{xy} dx dy \]

OpenStudy (anonymous):



OpenStudy (anonymous):

which triangle to use

OpenStudy (anonymous):

The bottom triangle because x goes from y to 1 which means FROM the line to the vertical line of x=1.

OpenStudy (anonymous):



OpenStudy (anonymous):

then i know my integrals change from 0 to x and from 0 to 1

OpenStudy (anonymous):

i dont understand then how to solve the integral

OpenStudy (anonymous):

So you have: \[\int\limits_0^1 \int\limits_0^y x^2 e^{xy} dy dx\]?

OpenStudy (anonymous):

First int should be 0 to x ****

OpenStudy (anonymous):

yes i have o to x

OpenStudy (anonymous):

i dunno how to solve x^2 e^(xy) i know its by parts but i am messing up

OpenStudy (anonymous):

Well if you integrate y first you don't need by parts: \[\int\limits_0^x x^2 (e^{x})^y dy = x^2 \ln(e^{x})e^{x y}|_0^x=x^3(e^{x^2}-1)\] From here you need to be clever: \[\int\limits_0^1 x^3(e^{x^2}-1)dx; \xi =x^2 \implies d \xi = 2x dx \implies dx = \frac{d \xi}{2x}\] From this we see the integral transforms into: \[\frac{1}{2}\int\limits_0^1 \xi(e^{\xi}-1)d \xi\] From here you can distribute xi and do integrate by parts on the first term and power rule on the second. Sorry it took so long to get back to you.